# Towards Global AI Inclusivity: A Large-Scale Multilingual Terminology Dataset

### v1

@@ Appendix C Additional Experiment Results
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 8.24 9.33 9.13
Bleu score std 10.82 11.97 11.65
Bleu 1gram precision mean 8.33 9.41 9.22
Bleu 1gram precision std 10.95 12.10 11.79
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 108.20 113.09 112.40
Bleu ref len std 45.16 43.78 44.13
Bleu pred len mean 112.33 117.01 116.40
Bleu pred len std 47.46 46.19 46.72
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.07 0.05 0.05
Bleu ratio mean 1.03 1.03 1.03
Bleu ratio std 0.10 0.08 0.09
Chrf score mean 8.61 9.72 9.53
Chrf score std 11.04 12.30 11.97
Ter score mean 101.41 100.10 100.54
Ter score std 16.93 16.77 16.69
Chrfpp score mean 8.61 9.72 9.53
Chrfpp score std 11.04 12.30 11.97
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 9.42 9.98 9.84
Bleu score std 11.49 12.78 12.60
Bleu 1gram precision mean 9.48 10.03 9.89
Bleu 1gram precision std 11.54 12.83 12.64
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 108.85 112.65 112.27
Bleu ref len std 45.86 43.78 44.25
Bleu pred len mean 113.57 117.39 117.19
Bleu pred len std 47.70 45.75 46.51
Bleu bp mean 0.99 1.00 0.99
Bleu bp std 0.05 0.03 0.04
Bleu ratio mean 1.05 1.04 1.04
Bleu ratio std 0.09 0.07 0.08
Chrf score mean 9.91 10.46 10.32
Chrf score std 11.91 13.19 12.99
Ter score mean 100.59 99.24 99.44
Ter score std 18.73 17.23 17.20
Chrfpp score mean 9.91 10.46 10.32
Chrfpp score std 11.91 13.19 12.99
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 10.55 10.60 10.58
Bleu score std 12.87 12.75 12.82
Bleu 1gram precision mean 10.58 10.63 10.61
Bleu 1gram precision std 12.99 12.85 12.91
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 116.51 119.64 120.03
Bleu ref len std 46.42 44.58 44.87
Bleu pred len mean 121.10 124.69 125.10
Bleu pred len std 48.28 46.61 46.94
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.04 1.04 1.04
Bleu ratio std 0.05 0.05 0.05
Chrf score mean 10.99 11.05 11.03
Chrf score std 13.18 13.04 13.11
Ter score mean 98.39 98.32 98.27
Ter score std 18.10 17.19 17.31
Chrfpp score mean 10.99 11.05 11.03
Chrfpp score std 13.18 13.04 13.11
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 9.22 9.45 9.57
Bleu score std 12.23 12.32 12.84
Bleu 1gram precision mean 9.25 9.48 9.60
Bleu 1gram precision std 12.35 12.44 12.95
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 114.14 115.01 114.24
Bleu ref len std 46.74 46.44 46.64
Bleu pred len mean 120.26 121.02 120.36
Bleu pred len std 49.41 48.96 49.29
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.02 0.02 0.02
Bleu ratio mean 1.06 1.06 1.06
Bleu ratio std 0.08 0.08 0.08
Chrf score mean 9.68 9.93 10.05
Chrf score std 12.48 12.60 13.09
Ter score mean 100.30 99.87 99.72
Ter score std 17.73 17.26 17.76
Chrfpp score mean 9.68 9.93 10.05
Chrfpp score std 12.48 12.60 13.09
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 8.11 8.08 8.16
Bleu score std 10.28 10.12 9.93
Bleu 1gram precision mean 8.12 8.09 8.17
Bleu 1gram precision std 10.29 10.13 9.94
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 119.56 119.59 119.77
Bleu ref len std 46.73 46.56 46.72
Bleu pred len mean 126.61 126.63 126.85
Bleu pred len std 48.80 48.64 48.80
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.06 1.06 1.06
Bleu ratio std 0.07 0.07 0.07
Chrf score mean 8.59 8.57 8.65
Chrf score std 10.70 10.55 10.35
Ter score mean 101.99 101.96 101.92
Ter score std 16.22 15.81 15.79
Chrfpp score mean 8.59 8.57 8.65
Chrfpp score std 10.70 10.55 10.35
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 19.18 19.07 18.94
Bleu score std 19.82 19.23 19.06
Bleu 1gram precision mean 19.24 19.13 19.00
Bleu 1gram precision std 19.82 19.25 19.07
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 195.84 220.60 200.93
Bleu ref len std 83.60 84.78 84.41
Bleu pred len mean 196.94 221.81 202.15
Bleu pred len std 84.71 85.63 85.45
Bleu bp mean 0.99 1.00 1.00
Bleu bp std 0.03 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 19.63 19.56 19.43
Chrf score std 19.96 19.44 19.28
Ter score mean 93.26 93.54 93.63
Ter score std 24.11 23.06 22.85
Chrfpp score mean 19.63 19.56 19.43
Chrfpp score std 19.96 19.44 19.28
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 18.43 19.38 18.78
Bleu score std 17.90 18.51 17.46
Bleu 1gram precision mean 18.48 19.44 18.84
Bleu 1gram precision std 17.92 18.54 17.49
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 222.40 231.72 229.59
Bleu ref len std 84.52 83.39 82.20
Bleu pred len mean 223.39 232.74 230.60
Bleu pred len std 85.50 84.23 83.10
Bleu bp mean 0.99 1.00 1.00
Bleu bp std 0.03 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 18.88 19.86 19.26
Chrf score std 18.04 18.68 17.62
Ter score mean 94.15 93.12 93.77
Ter score std 22.06 22.29 21.14
Chrfpp score mean 18.88 19.86 19.26
Chrfpp score std 18.04 18.68 17.62
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 18.02 19.54 18.43
Bleu score std 19.69 20.97 19.57
Bleu 1gram precision mean 18.13 19.62 18.51
Bleu 1gram precision std 19.72 21.00 19.60
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 230.39 236.14 236.37
Bleu ref len std 90.66 85.39 85.34
Bleu pred len mean 230.15 236.19 236.33
Bleu pred len std 90.89 85.82 85.72
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.04 0.03 0.03
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.02 0.02
Chrf score mean 18.44 19.95 18.83
Chrf score std 19.83 21.10 19.70
Ter score mean 94.40 92.58 93.89
Ter score std 23.67 24.65 23.35
Chrfpp score mean 18.44 19.95 18.83
Chrfpp score std 19.83 21.10 19.70
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 16.39 17.76 17.01
Bleu score std 17.61 18.41 18.37
Bleu 1gram precision mean 16.47 17.84 17.09
Bleu 1gram precision std 17.69 18.48 18.45
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 232.81 234.38 233.02
Bleu ref len std 88.09 86.95 88.03
Bleu pred len mean 232.90 234.61 233.12
Bleu pred len std 88.38 87.21 88.27
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.02 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 16.78 18.16 17.40
Chrf score std 17.75 18.54 18.50
Ter score mean 96.28 94.55 95.44
Ter score std 21.39 21.92 21.93
Chrfpp score mean 16.78 18.16 17.40
Chrfpp score std 17.75 18.54 18.50
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 13.84 14.06 14.17
Bleu score std 13.57 14.08 13.26
Bleu 1gram precision mean 13.89 14.10 14.21
Bleu 1gram precision std 13.61 14.12 13.31
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 238.17 238.01 238.42
Bleu ref len std 87.01 87.13 87.09
Bleu pred len mean 240.60 240.41 240.79
Bleu pred len std 86.86 87.01 86.90
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.01 1.01 1.01
Bleu ratio std 0.04 0.04 0.04
Chrf score mean 14.32 14.53 14.66
Chrf score std 13.83 14.31 13.54
Ter score mean 99.77 99.58 99.37
Ter score std 17.18 17.63 16.80
Chrfpp score mean 14.32 14.53 14.66
Chrfpp score std 13.83 14.31 13.54
@@
Direct Translation Prompting Refinement Word Alignment
@@
Arabic 0.81 0.83 0.81
Chinese 0.80 0.82 0.81
French 0.79 0.83 0.77
Japanese 0.83 0.85 0.84
Russian 0.80 0.83 0.78
@@
Arabic 0.82 0.83 0.82
Chinese 0.81 0.83 0.81
French 0.84 0.85 0.83
Japanese 0.84 0.85 0.84
Russian 0.83 0.85 0.81
@@
Arabic 0.84 0.84 0.83
Chinese 0.86 0.86 0.86
French 0.85 0.85 0.84
Japanese 0.88 0.88 0.88
Russian 0.85 0.85 0.83
@@
Arabic 0.82 0.83 0.81
Chinese 0.84 0.85 0.84
French 0.83 0.84 0.82
Japanese 0.84 0.85 0.84
Russian 0.83 0.84 0.81
@@
Arabic 0.81 0.81 0.80
Chinese 0.82 0.83 0.82
French 0.81 0.82 0.80
Japanese 0.85 0.86 0.86
Russian 0.83 0.83 0.81

### v3

@@ Appendix C Additional Experiment Results
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 8.24 9.33 9.13 9.45
Bleu score std 10.82 11.97 11.65 12.14
Bleu 1gram precision mean 8.33 9.41 9.22 9.52
Bleu 1gram precision std 10.95 12.10 11.79 12.26
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 108.20 113.09 112.40 112.49
Bleu ref len std 45.16 43.78 44.13 44.27
Bleu pred len mean 112.33 117.01 116.40 116.95
Bleu pred len std 47.46 46.19 46.72 46.64
Bleu bp mean 0.99 0.99 0.99 0.99
Bleu bp std 0.07 0.05 0.05 0.05
Bleu ratio mean 1.03 1.03 1.03 1.04
Bleu ratio std 0.10 0.08 0.09 0.09
Chrf score mean 8.61 9.72 9.53 9.88
Chrf score std 11.04 12.30 11.97 12.47
Ter score mean 101.41 100.10 100.54 100.13
Ter score std 16.93 16.77 16.69 17.03
Chrfpp score mean 8.61 9.72 9.53 9.88
Chrfpp score std 11.04 12.30 11.97 12.47
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 9.42 9.98 9.84 10.33
Bleu score std 11.49 12.78 12.60 13.04
Bleu 1gram precision mean 9.48 10.03 9.89 10.37
Bleu 1gram precision std 11.54 12.83 12.64 13.07
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 108.85 112.65 112.27 112.77
Bleu ref len std 45.86 43.78 44.25 44.27
Bleu pred len mean 113.57 117.39 117.19 117.78
Bleu pred len std 47.70 45.75 46.51 46.28
Bleu bp mean 0.99 1.00 0.99 1.00
Bleu bp std 0.05 0.03 0.04 0.02
Bleu ratio mean 1.05 1.04 1.04 1.05
Bleu ratio std 0.09 0.07 0.08 0.07
Chrf score mean 9.91 10.46 10.32 10.82
Chrf score std 11.91 13.19 12.99 13.44
Ter score mean 100.59 99.24 99.44 98.82
Ter score std 18.73 17.23 17.20 17.49
Chrfpp score mean 9.91 10.46 10.32 10.82
Chrfpp score std 11.91 13.19 12.99 13.44
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 10.55 10.60 10.58 10.55
Bleu score std 12.87 12.75 12.82 12.63
Bleu 1gram precision mean 10.58 10.63 10.61 10.58
Bleu 1gram precision std 12.99 12.85 12.91 12.72
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 116.51 119.64 120.03 119.68
Bleu ref len std 46.42 44.58 44.87 44.74
Bleu pred len mean 121.10 124.69 125.10 124.88
Bleu pred len std 48.28 46.61 46.94 46.73
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.04 1.04 1.04 1.04
Bleu ratio std 0.05 0.05 0.05 0.05
Chrf score mean 10.99 11.05 11.03 11.02
Chrf score std 13.18 13.04 13.11 12.93
Ter score mean 98.39 98.32 98.27 98.33
Ter score std 18.10 17.19 17.31 17.28
Chrfpp score mean 10.99 11.05 11.03 11.02
Chrfpp score std 13.18 13.04 13.11 12.93
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 9.22 9.45 9.57 9.66
Bleu score std 12.23 12.32 12.84 12.77
Bleu 1gram precision mean 9.25 9.48 9.60 9.70
Bleu 1gram precision std 12.35 12.44 12.95 12.88
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 114.14 115.01 114.24 114.45
Bleu ref len std 46.74 46.44 46.64 46.48
Bleu pred len mean 120.26 121.02 120.36 120.50
Bleu pred len std 49.41 48.96 49.29 49.07
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.02 0.02 0.02 0.02
Bleu ratio mean 1.06 1.06 1.06 1.06
Bleu ratio std 0.08 0.08 0.08 0.08
Chrf score mean 9.68 9.93 10.05 10.15
Chrf score std 12.48 12.60 13.09 13.06
Ter score mean 100.30 99.87 99.72 99.58
Ter score std 17.73 17.26 17.76 17.72
Chrfpp score mean 9.68 9.93 10.05 10.15
Chrfpp score std 12.48 12.60 13.09 13.06
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 8.11 8.08 8.16 8.42
Bleu score std 10.28 10.12 9.93 10.61
Bleu 1gram precision mean 8.12 8.09 8.17 8.43
Bleu 1gram precision std 10.29 10.13 9.94 10.63
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 119.56 119.59 119.77 119.71
Bleu ref len std 46.73 46.56 46.72 46.53
Bleu pred len mean 126.61 126.63 126.85 126.70
Bleu pred len std 48.80 48.64 48.80 48.68
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.06 1.06 1.06 1.06
Bleu ratio std 0.07 0.07 0.07 0.07
Chrf score mean 8.59 8.57 8.65 8.93
Chrf score std 10.70 10.55 10.35 11.05
Ter score mean 101.99 101.96 101.92 101.60
Ter score std 16.22 15.81 15.79 16.26
Chrfpp score mean 8.59 8.57 8.65 8.93
Chrfpp score std 10.70 10.55 10.35 11.05
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 19.18 19.07 18.94 20.62
Bleu score std 19.82 19.23 19.06 20.24
Bleu 1gram precision mean 19.24 19.13 19.00 20.67
Bleu 1gram precision std 19.82 19.25 19.07 20.26
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 195.84 220.60 200.93 205.87
Bleu ref len std 83.60 84.78 84.41 85.10
Bleu pred len mean 196.94 221.81 202.15 207.15
Bleu pred len std 84.71 85.63 85.45 86.09
Bleu bp mean 0.99 1.00 1.00 1.00
Bleu bp std 0.03 0.02 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00 1.01
Bleu ratio std 0.03 0.03 0.03 0.03
Chrf score mean 19.63 19.56 19.43 21.13
Chrf score std 19.96 19.44 19.28 20.45
Ter score mean 93.26 93.54 93.63 91.68
Ter score std 24.11 23.06 22.85 24.19
Chrfpp score mean 19.63 19.56 19.43 21.13
Chrfpp score std 19.96 19.44 19.28 20.45
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 18.43 19.38 18.78 20.13
Bleu score std 17.90 18.51 17.46 18.74
Bleu 1gram precision mean 18.48 19.44 18.84 20.18
Bleu 1gram precision std 17.92 18.54 17.49 18.77
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 222.40 231.72 229.59 229.97
Bleu ref len std 84.52 83.39 82.20 82.79
Bleu pred len mean 223.39 232.74 230.60 231.06
Bleu pred len std 85.50 84.23 83.10 83.63
Bleu bp mean 0.99 1.00 1.00 1.00
Bleu bp std 0.03 0.02 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03 0.03
Chrf score mean 18.88 19.86 19.26 20.62
Chrf score std 18.04 18.68 17.62 18.91
Ter score mean 94.15 93.12 93.77 92.12
Ter score std 22.06 22.29 21.14 22.29
Chrfpp score mean 18.88 19.86 19.26 20.62
Chrfpp score std 18.04 18.68 17.62 18.91
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 18.02 19.54 18.43 19.62
Bleu score std 19.69 20.97 19.57 20.51
Bleu 1gram precision mean 18.13 19.62 18.51 19.70
Bleu 1gram precision std 19.72 21.00 19.60 20.54
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 230.39 236.14 236.37 235.79
Bleu ref len std 90.66 85.39 85.34 85.01
Bleu pred len mean 230.15 236.19 236.33 235.92
Bleu pred len std 90.89 85.82 85.72 85.40
Bleu bp mean 0.99 0.99 0.99 0.99
Bleu bp std 0.04 0.03 0.03 0.03
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.03 0.02 0.02 0.03
Chrf score mean 18.44 19.95 18.83 20.05
Chrf score std 19.83 21.10 19.70 20.66
Ter score mean 94.40 92.58 93.89 92.47
Ter score std 23.67 24.65 23.35 24.12
Chrfpp score mean 18.44 19.95 18.83 20.05
Chrfpp score std 19.83 21.10 19.70 20.66
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 16.39 17.76 17.01 17.75
Bleu score std 17.61 18.41 18.37 18.79
Bleu 1gram precision mean 16.47 17.84 17.09 17.82
Bleu 1gram precision std 17.69 18.48 18.45 18.86
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 232.81 234.38 233.02 233.86
Bleu ref len std 88.09 86.95 88.03 86.89
Bleu pred len mean 232.90 234.61 233.12 234.16
Bleu pred len std 88.38 87.21 88.27 87.17
Bleu bp mean 0.99 0.99 0.99 0.99
Bleu bp std 0.02 0.02 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03 0.03
Chrf score mean 16.78 18.16 17.40 18.15
Chrf score std 17.75 18.54 18.50 18.94
Ter score mean 96.28 94.55 95.44 94.59
Ter score std 21.39 21.92 21.93 22.42
Chrfpp score mean 16.78 18.16 17.40 18.15
Chrfpp score std 17.75 18.54 18.50 18.94
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 13.84 14.06 14.17 14.53
Bleu score std 13.57 14.08 13.26 14.74
Bleu 1gram precision mean 13.89 14.10 14.21 14.57
Bleu 1gram precision std 13.61 14.12 13.31 14.77
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 238.17 238.01 238.42 238.06
Bleu ref len std 87.01 87.13 87.09 87.13
Bleu pred len mean 240.60 240.41 240.79 240.49
Bleu pred len std 86.86 87.01 86.90 86.95
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.01 1.01 1.01 1.01
Bleu ratio std 0.04 0.04 0.04 0.04
Chrf score mean 14.32 14.53 14.66 15.02
Chrf score std 13.83 14.31 13.54 15.01
Ter score mean 99.77 99.58 99.37 98.98
Ter score std 17.18 17.63 16.80 18.37
Chrfpp score mean 14.32 14.53 14.66 15.02
Chrfpp score std 13.83 14.31 13.54 15.01
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
@@
Arabic 0.81 0.83 0.81 0.83
Chinese 0.80 0.82 0.81 0.83
French 0.79 0.83 0.77 0.80
Japanese 0.83 0.85 0.84 0.86
Russian 0.80 0.83 0.78 0.82
@@
Arabic 0.82 0.83 0.82 0.84
Chinese 0.81 0.83 0.81 0.84
French 0.84 0.85 0.83 0.85
Japanese 0.84 0.85 0.84 0.86
Russian 0.83 0.85 0.81 0.85
@@
Arabic 0.84 0.84 0.83 0.84
Chinese 0.86 0.86 0.86 0.86
French 0.85 0.85 0.84 0.85
Japanese 0.88 0.88 0.88 0.88
Russian 0.85 0.85 0.83 0.85
@@
Arabic 0.82 0.83 0.81 0.83
Chinese 0.84 0.85 0.84 0.85
French 0.83 0.84 0.82 0.84
Japanese 0.84 0.85 0.84 0.86
Russian 0.83 0.84 0.81 0.84
@@
Arabic 0.81 0.81 0.80 0.82
Chinese 0.82 0.83 0.82 0.83
French 0.81 0.82 0.80 0.82
Japanese 0.85 0.86 0.86 0.86
Russian 0.83 0.83 0.81 0.84
@@
Figure 13: Arabic translation examples of the two integration cases on nllb, including one good example and one bad example for each case.
Figure 14: Chinese translation examples of the two integration cases on nllb, including one good example and one bad example for each case.
Figure 15: French translation examples of the two integration cases on nllb, including one good example and one bad example for each case.
Figure 16: Japanese translation examples of the two integration cases on nllb, including one good example and one bad example for each case.
Figure 17: Russian translation examples of the two integration cases on nllb, including one good example and one bad example for each case.